OpenStudy (babynini):

Write z1 and z2 in polar form, and then find z1(z2) and z1/z2. Rewrite the results in the form a+bi

OpenStudy (babynini):

z1= -5i z2=2+21

OpenStudy (babynini):

@zepdrix :)

OpenStudy (babynini):

for z1 polar I got = 5cis(3pi/2) for z2 polar I got = 2(sq2)cis(5pi4)

OpenStudy (babynini):

Is that correct?

OpenStudy (perl):

z2 is incorrect

OpenStudy (perl):

z2 = 2 + 2i ?

OpenStudy (babynini):

yeah, hm

OpenStudy (perl):

z2= 2(sq2)cis(pi4)

OpenStudy (babynini):

Doesn't it have to be where theta is positive? ie. Quadrant 3?

OpenStudy (perl):

theta is positive in quadrant 1

OpenStudy (perl):

im not sure how you got that angle, the complex number 2 + 2i is in quadrant 1 , in the complex plane

OpenStudy (babynini):

oh shoot. Yeah, i see now.

OpenStudy (babynini):

fail. haha thanks.

OpenStudy (babynini):

ok, three parts left :)

OpenStudy (babynini):

z1 times z2

OpenStudy (babynini):

let me try it out on paper.

OpenStudy (babynini):

z1(z2)= 10(sq2)cis(7pi/4)

OpenStudy (perl):

\[\large \rm if ~ z_1 =r_1\angle \theta_1,~ z_2 =r_2\angle \theta_2 \\ \large z_1 \cdot z_2 = r_1\cdot r_2~ \angle \left ( \theta_1 + \theta_2 \right )\]

OpenStudy (perl):

yes thats correct

OpenStudy (babynini):

what i'm working with is: \[z _{1}z _{2}=r _{1}r _{2}(\cos(\theta _{1}+\theta _{2})+isin(\theta _{1}+\theta _{2}))\]

OpenStudy (babynini):

yay! next is z1/z2

OpenStudy (perl):

divide radii, subtract angles

OpenStudy (babynini):

ok:) let me write it out.

OpenStudy (babynini):

er is 5/2(sq2) = - [5(sq10)]/4?

OpenStudy (perl):

5/4 sqrt(2)

OpenStudy (babynini):

the root cannot be in the base :o

OpenStudy (perl):